# Moduli spaces and breather dynamics of analytic solutions in chiral magnets without Heisenberg exchange interaction

### v1

@@ 4 Static solutions in 3D
However, it will not be possible to assemble a Hopfion solution, since it is a necessary condition that preimages of nearly all values of $n_{3}$ should form closed loops. The preimages of $n_{3}$ in any solution we assemble in the way we described above, will have no values of $n_{3}$ for which the preimages form a loop. A Hopfion may exist dynamically, however.
@@ 5 Conclusion
In this paper we consider the limit of the magnetic Skyrme theory – the theory of magnetic textures in chiral magnetic materials, where the Heisenberg exchange interaction is absent or negligible. Formulating the equations in terms of a fluid-flow equation, we find the interesting result that an enlarged moduli space emerges. The fluid lines are all circles, but due to the missing Heisenberg energy term, the circles can be moved at will – as long as they do not intersect one another. This leads to the interpretation that the moduli space is that of a “string” with the speed limit, corresponding to the circles being non-intersecting. We call the maps that generate the moduli space “contour-preserving” maps. We give a few explicit examples of exact solutions that are constructed in this way and confirm that they obey the equations of motion for the fields. We further go on and contemplate the case of time-dependent solutions that correspond to breathers. Finally, we conclude that no knotted solitons or Hopfions exist in this model.
@@ Acknowledgements
Appendix A Sanity check for fluid flow equations
Let us check that these solutions are valid in a more direct way. Firstly, the static Euler-Lagrange equations in the case of the axisymmetric potential (7) reduce to
$\bm{n}\cdot\nabla n_{3}=0\ ,$
$2k\bm{n}\cdot\nabla\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ .$ (A1)
Let us consider one of these curves of constant $n_{3}$. We can give this a unit-speed parametrization respecting the orientation given by the in-plane vector, $\vec{X}(\sigma)$, i.e. $\lvert\vec{X}^{\prime}(\sigma)\rvert=1$, and $\vec{X}^{\prime}(\sigma)\parallel\bm{n}(\vec{X})$. We can then introduce the Frenet frame: $\vec{T}(\sigma)=\vec{X}^{\prime}(\sigma)$, $\vec{N}=\vec{T}^{\prime}(\sigma)/\lvert\vec{T}^{\prime}(\sigma)\rvert$ form an orthonormal pair of vectors at each point on the curve $\vec{X}$, with $\vec{T}$ tangent to the curve and $\vec{N}$ normal. The signed curvature $\kappa$ is given by the Frenet-Serret equations:
$\frac{\mathop{}\!\mathrm{d}\vec{T}}{\mathop{}\!\mathrm{d}\sigma}=\kappa\vec{N}\ ,\quad\quad\frac{\mathop{}\!\mathrm{d}\vec{N}}{\mathop{}\!\mathrm{d}\sigma}=-\kappa\vec{T}\ .$ (A2)
Note that we can only do this because $\kappa$ never vanishes for the curves in question. However, since we do have a Frenet frame, we can create a local co-ordinate system $(\sigma,\tau)$ in the vicinity of the curve, $\vec{x}=\vec{X}(\sigma)+\tau\vec{N}(\sigma)$. We can then transform the static Euler-Lagrange equations to this co-ordinate system:
$\underbrace{\bm{n}\cdot\vec{T}\partial_{\sigma}n_{3}}_{\text{vanishes}}+\bm{n}\cdot\vec{N}\partial_{\tau}n_{3}=0\ ,$ (A3)
$2k\bm{n}\cdot\vec{T}\partial_{\sigma}\nu+2k\bm{n}\cdot\vec{N}\partial_{\tau}\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ .$ (A4)
Now we substitute in our solution (23), where curves of constant $n_{3}$ are circles with $\vec{\nu}$ tangent to those curves with signed radius $R(n_{3})$ given by the formula (12). (So $\kappa$ is constant and equal to $1/R(n_{3})$.) Since $\vec{\nu}$ is tangent to the curve, the first equation is trivially satisfied, then for the second equation we apply the Frenet-Serret equations:
$\partial_{\sigma}\nu=\kappa\mathrm{i}\nu=\frac{\mathrm{i}}{R(n_{3})}\nu=\frac{-\mathrm{i}f^{\prime}(n_{3})}{2k\sqrt{1-n_{3}^{2}}}\nu=\frac{-\mathrm{i}f^{\prime}(n_{3})}{2k\bm{n}\cdot\vec{T}}\nu\ ,$ (A5)
and we see that it is satisfied also.
In general, the locally adapted co-ordinates of the Frenet frame cannot be extended to global ones. However, we have shown the static Euler-Lagrange equations are satisfied locally everywhere so this is enough for a formal validity check.
The main point is that the equation (A1) depends on the derivatives only through the combination $\bm{n}\cdot\nabla=\nu\cdot\nabla$. So, once a specific solution is found, the lines tangent to $\nu$ (circles in our case), are locally solutions irrespective to what happens in their neighbourhood. Thus the lines tangent to $\nu$ can be displaced with different translation vectors without spoiling the validity of the equation of motion.
Appendix B Radially symmetric restricted solutions
Considering the potential (7) and reparametrizing it by
$h=\tilde{h}\sin\alpha\ ,\qquad u=\frac{\tilde{h}}{2}\cos\alpha\ .$ (A6)
we have that
$-f^{\prime}(n_{3})=\tilde{h}(\sin\alpha+n_{3}\cos\alpha)\ .$ (A7)
It is important that the potential only depends on $n_{3}$, otherwise further symmetry breaking will occur. For $40\leq\alpha<\pi/$ 4 we have a metastable point at $n_{3}=-1$. For $3\pi/4<\alpha<\pi$ we have spontaneous symmetry breaking, and $\alpha=3\pi/4$ is the line of the $(1-n_{3})^{2}$ potential that was considered in Ref. [21]. $\alpha=\pi/2$ corresponds to the Zeeman potential. We can therefore limit ourselves, without loss of generality, to $0\leq\alpha\leq\pi$, and $\tilde{h}\geq 0$.
We consider now the simplified case where $X_{c}:=0$, so that the equation of motion is simply given by Eq. (12) with $R=r$ being the radius in the 2-plane $r=\sqrt{x^{2}+y^{2}}$. Using $-f^{\prime}(n_{3})$ of Eq. (A7), we obtain
$r=\frac{2\ell\sin\Theta}{\sin\alpha+\cos\alpha\cos\Theta}\ ,$ (A8)
[…]
$\alpha=\pi/2\,:\quad\Theta(r)=\pi-\arcsin\rho(r)\ ,$ $\rho(r)$ $=\frac{r}{2\ell}\ ,$ (A9)
$\alpha=3\pi/4\,:\quad\Theta(r)=2\arctan\frac{1}{\rho(r)}\ ,$ $\rho(r)$ $=\frac{r}{2\sqrt{2}\ell}\ ,$ (A10)
with $\rho=\rho(r)\in[0,1)$. In general what happens is the following. For $40\leq\alpha\leq\pi/$ 4, the solutions exist on the entire range $\rho\in[0,\infty]$, but none of them reach $\Theta(\infty)\to 0$ and as a result they are fraction solutions with topological charges $Q\leq\frac{1}{2}$.
For $\pi/4<\alpha<3\pi/4$ the situation is similar to $\alpha=\pi/2$, we have a supercompacton as restricted solution, but again the topological charge is $10<Q<$ 1. As $\alpha\to 3\pi/4$ the supercompacton becomes a lump and the topological charge becomes integer, $Q=1$. The generic solution for $0\leq\alpha\leq\pi$ is
$\Theta(r)=\arccos\left(-\frac{2\sqrt{4+\rho^{2}\cos 2\alpha}+\frac{1}{2}\rho^{2}\sin 2\alpha}{4+\rho^{2}\cos^{2}\alpha}\right),\qquad\rho=\frac{r}{\ell}\ ,$ (A11)
with
$\rho\in\begin{cases}[0,\infty]\ ,&0\leq\alpha\leq\frac{\pi}{4}\ ,\\ \left[0,2\sqrt{\frac{-1}{\cos 2\alpha}}\right]\ ,&\frac{\pi}{4}<\alpha<\frac{3\pi}{4}\ ,\\ [0,\infty]\ ,&\frac{3\pi}{4}\leq\alpha\leq\pi\ .\end{cases}$ (A12)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Restricted solutions of magnetic Skyrmions with the potential (7), parametrized by Eq. (A6) with $\alpha$ in the range $0\leq\alpha\leq\pi$.
In the range $3\pi/4<\alpha\leq\pi$ we have again a fracton. The fractons that exist for $40\leq\alpha\leq\pi/$ 4 and for $3\pi/4<\alpha\leq\pi$ exist on the entire range $\rho\in[0,\infty]$ and have the asymptotic behaviour
$\lim_{r\to\infty}\Theta(r)=\arccos(-\tan\alpha)\ ,$ (A13)
which is the true vacuum of the potential. Notice that as soon as we introduce the $E_{2}$ term, this solution acquires a “log” divergent energy due to the angular derivative. So it is like a global vortex solution. The fracton has a definite topological charge and Eq. (2) reduces to
$-\frac{1}{2}\int_{\pi}^{\Theta(\infty)}\sin(\Theta)\mathop{}\!\mathrm{d}\Theta=-\frac{1}{2}[\cos\Theta]_{\Theta(\infty)}^{\pi}=\frac{1}{2}(1+\tan\alpha)\ ,\qquad\alpha\in\big{\{}[0,\pi/4],(3\pi/4,\pi]\big{\}}\ ,$ (A14)
which becomes $1/2$ for $\alpha=0,\pi$; in this case we can call it a meron.
Appendix C The restricted model in complex co-ordinates
@@
=±$\mathcal{E}=\pm\mathrm{i}k\frac{2\bar{\partial}\bar{\nu}-2\partial\nu+\bar{\nu}^{2}\bar{\partial}\nu-\nu^{2}\partial\bar{\nu}+\nu\bar{\nu}(-\bar{\partial}\bar{\nu}+\partial\nu)}{2\sqrt{1-\nu\bar{\nu}}}+f(\pm\sqrt{1-\nu\bar{\nu}})\ ,$ (A16)
@@
$\left(2\mathrm{i}\nu\partial+2\mathrm{i}\bar{\nu}\bar{\partial}-\frac{f^{\prime}}{k}\right)\nu=0\ ,$ (A17)
which is indeed very elegant and independent of the sign $\pm$ choosing the patch. A simple solution that covers the north patch, suitable for the Zeeman potential which has $f^{\prime}=-h$, is given by

### v2

@@ 4 Static solutions in 3D
However, it will not be possible to assemble a static Hopfion solution: a Hopfion by definition exists within a uniform background $\bm{n}=\bm{N}$ some minimum of $V$, and its topology is defined with respect to that. The Hopfion attains all values of $\bm{n}$ in the interior of its domain. In particular, it attains a value of $\bm{n}$ which is a local maximum of $V$. If a Hopfion existed in this model, if it attained this $\bm{n}$ at any point it would have to attain it along a straight line through that point, which would intersect with the uniform background, giving a contradiction. A Hopfion may exist dynamically, however.
@@ 5 Conclusion
In this paper we consider the limit of the magnetic Skyrme theory – the theory of magnetic textures in chiral magnetic materials, where the Heisenberg exchange interaction is absent or negligible. Formulating the equations in terms of a fluid-flow equation, we find the interesting result that an enlarged moduli space emerges. The fluid lines are all circles, but due to the missing Heisenberg energy term, the circles can be moved at will – as long as they do not intersect one another. This leads to the interpretation that the moduli space is that of a “string” with the speed limit, corresponding to the circles being non-intersecting. We give a few explicit examples of exact solutions that are constructed in this way and confirm that they obey the equations of motion for the fields. We further go on and contemplate the case of time-dependent solutions that correspond to breathers. Finally, we conclude that no knotted solitons or Hopfions exist in this model.
@@
All conclusions in this paper followed from looking for continuous solutions to the equations of motion, and thus often found configurations only on a finite disc. More generally, if we allow discontinuous weak solutions, we would expect to find solutions with the whole plane as a domain, but we must deal with the behaviour of the boundary where $\bm{n}$ is discontinuous, using the theory of shock and rarefaction waves. Since the shock condition is trivially satisfied if the fluid flow on both sides of the shock are parallel to the shock boundary, we can look for continuous solutions on subsets of the plane with boundary conditions that $\vec{\nu}$ is parallel to the boundary. These can be assembled to create discontinuous solutions in the whole plane, but conversely we cannot expect to find all discontinuous solutions in the plane by assembling these particular ‘self-contained’ continuous solutions.
@@ Acknowledgements
Appendix A The restricted model in complex co-ordinates
@@
$\mathcal{E}=\pm\mathrm{i}k\frac{2\bar{\partial}\bar{\nu}-2\partial\nu+\bar{\nu}^{2}\bar{\partial}\nu-\nu^{2}\partial\bar{\nu}+\nu\bar{\nu}(-\bar{\partial}\bar{\nu}+\partial\nu)}{2\sqrt{1-\nu\bar{\nu}}}+U(\pm\sqrt{1-\nu\bar{\nu}})\ ,$ (A2)
@@
$\left(2\mathrm{i}\nu\partial+2\mathrm{i}\bar{\nu}\bar{\partial}-\frac{U^{\prime}}{k}\right)\nu=0\ ,$ (A3)
which is indeed very elegant and independent of the sign $\pm$ choosing the patch. A simple solution that covers the north patch, suitable for the Zeeman potential which has $U^{\prime}=-h$, is given by